ABC Company has the following balances at the end of their fiscal year: Cash 11,000 Accounts Receivable 3,000 Long Term Debt 4,000 Accounts Payable 2,000 Inventory 2,000

Based on only these account balances, calculate the total owners equity at the end of the year. ( (Answer in whole $s without the $ sign or comma, xxxxx, e.g. 12000.)
Answer #1

Total assets=Cash+AR+Inventory

=(11000+3000+2000)=$16,000

Total liabilities=Long term debt+AP

=(4000+2000)=$6000

Total assets=Total liabilities+total owners equity

Total owners equity=$16,000-$6000

=10000